DIODE LASER SPECTROSCOPY MEASUREMENT OF OXYGEN A-BAND LINESTRENGTHS

\underline{ROBERT J. BRECHA}, Physics Dept., Univ. of Dayton, Dayton, OH 45469-2314.

Linestrength measurements for the oxygen A-band at 13122 cm^{−1} using tunable diode laser absorption spectroscopy are reported. The diode lasers are scanned over individual lines and the absorption at different pressures is recorded; from these data the linestrength for each transition is obtained. Average values for each transition are compared to those previously published, and in particular a comparison with the HITRAN database will be presented. Preliminary results indicate linestrengths 5-10% lower than HITRAN values, but in approximate agreement with another set of published results.